OpenStudy (anonymous):

find y'' if y=8xsinx I am getting y''=8xsinx+16cosx But that is not one of the choices What am I doing wrong?

OpenStudy (johnweldon1993):

what are the choices first?

OpenStudy (nubeer):

u have to find double derivative right? what is the first derivative u got?

OpenStudy (nubeer):

and yeah this looks wrong.. there should be 1 more term i think

OpenStudy (anonymous):

\[A. y''=8cosx-16xsinx\] \[B. y''=16cosx-8xsinx\] \[C. y''=-16cosx+8xsinx\] \[D. y''=-8xsinx\] and yes I have to find the double derivative

OpenStudy (anonymous):

well I got y=8xsinx then y' should be \[y'=8x \frac{ d }{ dx }sinx+sinx \frac{ d }{ dx }(8x)\] \[y'=8xcosx+8sinx\]

OpenStudy (nubeer):

yes this is right..

OpenStudy (johnweldon1993):

well this looks like it can have the product rule used doesn't it let f be 8x and let g be sin x you do f' times g + g' times f so 8 times sinx = 8sinx as i typed...you came to the correct way lol :)

OpenStudy (anonymous):

so then I tried taking the derivative of y' \[y''=(8x \frac{ d }{ dx }cosx+cosx \frac{ d }{ dx }(8x))+(8\frac{ d }{ dx }sinx+sinx \frac{ d }{ dx}(8))\] which should equal \[y''=8xsinx+cosx(8)+8cosx\]

OpenStudy (nubeer):

nop.. derivative of cosx is -sinx.. thats the mistake ..

OpenStudy (johnweldon1993):

there we go ...as @nubeer has said

OpenStudy (anonymous):

ok so it is y''=-8xsinx+16cosx which is choice B right?

OpenStudy (nubeer):

yes..

OpenStudy (johnweldon1993):

yes it is

OpenStudy (anonymous):

thanks guys

OpenStudy (johnweldon1993):

your welcome :)

OpenStudy (nubeer):

anytime..

OpenStudy (anonymous):

I am still trying to figure out a volume of a rectangle one

OpenStudy (anonymous):

word problem volume of a rectangular box with a square base remains constant at 1000cm^3 as the area of the base increases at a rate of 2cm^2/sec. Find the rate at which the height of the box is decreasing when each side of the base is 19cm long

OpenStudy (anonymous):



OpenStudy (anonymous):

so if A*h=1000 and the given rate of change is 2cm^2/sec then do I \[A*\frac{ d }{ dt }(h)+h*\frac{ d }{ dt }(A)=\frac{ d }{ dt }1000\]

OpenStudy (nubeer):

hmm no clue on this one sorry

OpenStudy (anonymous):

thanks anyway